OpenStudy (anonymous):

If a species contains 23% A in its DNA what is the percentage of guanine it would contain?

OpenStudy (shrutipande9):

purine to pyrimidine ratio should be 1:1 also A pairs with T and G with C. so % of A=% of T % of G=% of C\[\frac{ A+G }{ C+T }=\frac{ 1 }{ 1 }\]

OpenStudy (shrutipande9):

i hope u can solve this further...if u have a problem..tag me:)